# ToDo (#1): #2

### v2

@@ 2.2.3 Beam Dynamics
Beam simulations for the LEBT have been performed using IBSimu and Warp. IBSimu provides a more accurate plasma model, but is computationally intensive and so was used only for the region by the ion source where the plasma is extracted and the density is thus high, requiring this accuracy. IBSimu is used for the first several mm after the extraction aperture where density remains high. After this, and for the remainder of the downstream region of the LEBT, the distributions from IBSimu are handed off to Warp which provides accurate results for the lower density regions of the beam with less computational burden.
For the IBSimu simulations a 12 mA beam composed of 80% H+${}_{2}^{+}$, 10% ${}^{+}$, and 10% H+${}_{3}^{+}$ was used. This provides ample current for IsoDAR assuming 50% global transmission efficiency while using current from the ion source that has already been empirically demonstrated. As illustrated in Fig. 2.6 and in Table 2.2 the LEBT is able to match the required input parameters for the RFQ. Moreover, as illustrated in Fig. 2.7 this match is achieved with emittance that surpasses that used for baseline RFQ simulations.
Figure 2.6: A temperature plot showing the Twiss parameters required for optimal transmission through the RFQ. The indicated point shows the parameters at the end of the LEBT.
Figure 2.7: Phase space from the LEBT simulation (left) vs phase space for ideal beam input to the RFQ (right).
@@ 2.3 The Radiofrequency Quadrupole
Figure 2.8: Cut view of the IsoDAR RFQ. In the split-coaxial design, vanes are attached to the front and end plates rather than radially to the mantle.
@@ 2.3.1 Technical Design
The technical design was presented in Refs. [35, 36, 37] and the optimization of the RF loop coupler in Ref. [38]. The thermal properties and cooling considerations were presented in Ref. [39]. An optimization scheme using machine learning was presented in Ref. [40].
Figure 2.9: Modes of the mechanical deformation of the RFQ vanes. The frequencies are 18 Hz and 25 Hz.
@@ 2.3.2 Beam Dynamics
Beam dynamics studies were performed of the full 3D particle distribution resulting from the LEBT simulations going through the RFQ, including the effects of the entrance gap and asymmetry (due to the split coaxial design) and the exit gap and asymmetry. Fine tuning of the phase of the exit gap (i.e. tuning the length of a drift preceding the exit gap), leads to final output beam parameters listed in Tab. 2.4.
@@
Phase width (1-RMS) ${}^{\circ}$
@@
Figure 3.1: Isometric view of the cyclotron
In this section we describe the details of the cyclotron design at the level of a preliminary design report.
@@ 3.1 Magnet Yoke
The magnet design presented here will necessarily be modified for a final design, as the specific details of the extraction orbit and all required extraction elements have not been finalized.
The engineering design of the IsoDAR magnet benefits from the experience gained by IBA in the fabrication of its C400 cyclotron, which is slightly larger, at 7 meters outer diameter.
Figure 3.2: C400 on the milling machine, C230 in front of it.
@@ 3.1.1 Design and fabrication of the cyclotron magnet
The cyclotron is divided in the same fashion as the C400, into smaller parts facilitating transport to the site. The magnet is split radially and will be assembled on site. The vacuum chamber is split in 3 (self-centering) parts along its height, the vacuum tightness is ensured by 2 axial O-rings at each interface except for the interface between the upper and lower chambers were 2 radial O-rings are used. An exploded of the upper and lower parts is shown in Figs. 3.4 and 3.4. The materials used for the magnet parts is cyclotron steel (the exact composition shall be determined after a detailed physics study), the vacuum chambers will be made of stainless steel.
Figure 3.3: Cyclotron upper steel exploded.
@@
An overview of a cut in the cyclotron is shown in Fig. 3.5. The external and vacuum disks are connected to each other as well as to the flux return using specific bolts (such as the Superbolt from Nord-Lock). The poles are split in several parts and an easily removable part should be added in the next iteration to ease the mapping and possibly improve the magnetic performance. A detailed view of the cyclotron components is shown in Fig. 3.6.
@@
Table 3.1 gives some more information regarding the yoke parts.
@@ 3.2 Interface with the RFQ
Figure 3.7: RFQ overview
The RFQ will be outside of the cyclotron’s vacuum, therefore a vacuum sleeve will be used to ensure the integrity of the cyclotron’s vacuum. This sleeve will be fixed on top of the upper external disk and the vacuum sealing will be done radially on the upper internal disk (as seen in figure number 3.9 ). The O-rings could have a cross section diameter of 8mm and an inner diameter of 305mm.
In order to allow some adjustement, the RFQ sleeve allows a rotation of 5mrad around the lower O-ring contact patch (discussed with the manufacturer of the RFQ).
Figure 3.8: RFQ zoom
The RFQ sleeve is as thin as possible to avoid any loss in magnetic field in the central region (hence the O-rings in the Yoke parts).
Figure 3.9: RFQ cut section 1
Fig. 3.10 shows that the vacuum sealing will be done at the bottom of the RFQ using a thick O-ring (diameter 10 mm) which will be compressed by the RFQ’s weight ($\approx 3000$ N). The dimensions of the vacuum sleeve are listed in Tab. 3.2.
The adjustment feet shown in Fig. 3.8 are fixed to the cyclotron yoke and allow the following alignment modifications through four sets of four set screws one and vertical adjustment nut:
  • Rotation: ±5plus-or-minus5$\pm$ 5± ${}^{\circ}$
@@ 3.3 Central Region
Figure 3.11: CAD rendering of the spiral inflector and particle trajectories for early tests at Best Cyclotron Systems Inc. [41]. The positive and negative electrostatic deflectors are shown in blue and green respectively. The cyclotron magnetic field in this image is directed vertically upwards. Particles enter the spiral inflector via the rectangular grounded collimator (solid gray) and are guided into the cyclotron mid-plane by means of the cyclotron magnetic field and the electrostatic potential between the electrodes. The copper housing (transparent gray) isolates the spiral inflector from the RF fields driving the cyclotron.
@@
The injection system design is a crucial component to achieve the highest beam current. Our design balances injecting the $\mathrm{H}_{2}^{+}$ molecule at high energy against avoiding the use of complex high-voltage platforms. Higher energy injection mitigates space charge and to reduces geometrical (non-normalized) emittance. An energy of 35 keV/amu was selected which determined the output energy of the RFQ. As in commercial cyclotrons, an axial injection system based on a Spiral Inflector (SI) will be used to bend the beam from the axial direction to the median plane. This concept is visualized in Figure 3.11 using trajectories calculated with OPAL [42]. The exit point of the spiral inflector is also the starting point of acceleration in the Central Region (CR) of the cyclotron. The central region and the spiral inflector have to be designed carefully and together because they have strong interplay. This procedure has already been performed and tested for the injection tests during the summers of 2013 and 2014 at the Best Cyclotron Systems Inc. (BCSI) laboratory in Vancouver. In the following subsection, we will describe the design process of the central region for the BCS test stand cyclotron followed by a preliminary design for the full IsoDAR machine.
3.3.0.1 Spiral Inflector Design
The axial injection of the ionized beam into a cyclotron is realized using an electrostatic device called a spiral inflector, which consists of two curved electrode deflectors (see Figure 3.11). The electrostatic potential between these electrodes is able to bend the beam ${}^{\circ}$ from the axial line to the median plane of the cyclotron. The helical trajectory of the beam is determined both by the shape of the electrodes (electric field) and the magnetic field of the cyclotron as the beam is bent into the median plane. The applied voltage depends on the velocity of the beam as well as the rigidity of the ions.
Figure 3.12: Photograph of the spiral inflector used during the tests at Best Cyclotron Systems.
The spiral inflector tested at BCSI (shown in the photograph in Figure 3.12) was designed to mimic the conditions required for the IsoDAR inflector, its primary defining characteristic being the 15 mm gap between the electrodes. Compared to other similar designs, this is rather large because it has to take into account the larger beam size due to space-charge effects and the higher rigidity of $\mathrm{H}_{2}^{+}$. Due to the large distance between the electrodes, the SI occupies a volume in which the magnetic field components’ variation is not negligible, as it is assumed in the analytical treatment of the spiral inflector. This effect has to be carefully taken into account in order to shape correctly the spiral inflector electrodes and to avoid the introduction of a high energy-spread during the beam transport through the device. The spiral inflector has a copper housing that surrounds the electrodes to minimize the interaction between the electric fields generated by the dees and the electrostatic fields between the electrodes. To minimize beam striking the uncooled electrodes, a water-cooled, grounded, rectangular collimator shields the spiral inflector entrance.
The preliminary design of the spiral inflector called for a nominal voltage of ±$\pm 11$ kV and a tilt angle of ${}^{\circ}$ for optimal beam injection at 60 keV. The total height of the device was approximately 80 mm. These parameters were found by using a MATLAB code based on an analytic theory for spiral inflectors. The electrode shape was calculated using VectorFields OPERA. This design was later modified to account for some effects due to fringe fields on the electrodes. One of these modifications was to reduce the overall length of each electrode by 5 mm.
@@ 3.3.0.1 Spiral Inflector Design
First turn acceleration is achieved by the dee tips extending into recesses in the spiral inflector housing. The shapes of the dee tips and the spiral inflector housing have been designed to guide the particles from the spiral inflector exit to the acceleration region while providing the necessary energy gain and beam focusing.
[…]
The main differences between the BCSI test stand design and the IsoDAR design are:
  • The IsoDAR cyclotron will be equipped with 4 accelerating RF cavities, while the test cyclotron used only 2 cavities.
  • The harmonic mode (and the frequency) will be decreased from 6th harmonic (49.2 MHz) to 4th harmonic (32.8 MHz).
  • The injection energy will be increased from 30 keV/amu to 35 keV/amu.
  • The spiral inflector gap will be increased from 15 mm to 16.5 mm to accommodate a larger beam.

### v3

@@ 2.2.3 Beam Dynamics
Beam simulations for the LEBT have been performed using IBSimu and Warp. IBSimu provides a more accurate plasma model, but is computationally intensive and so was used only for the region by the ion source where the plasma is extracted and the density is thus high, requiring this accuracy. IBSimu is used for the first several millimeters after the extraction aperture, where density remains high. After this, and for the remainder of the downstream region of the LEBT, the distributions from IBSimu are handed off to Warp which provides accurate results for the lower density regions of the beam with less computational burden.
Space charge is included in all calculations. However, as is well-known, in the absence of external electrical fields, slow secondary charged particles from the interaction of the beam with the residual gas can be accumulated inside the beam envelope, effectively neutralizing the beam charge. This process is called “space charge compensation” [winklehnerSpaceChargeCompensation2015]. In all LEBT simulations, we estimate space charge compensation according to the model presented in Ref. [winklehnerSpaceChargeCompensation2015]. It is usually above 80%, alleviating the detrimental effect of space charge in some parts of the LEBT. However, we also test our designs with no compensation and find that, while beam losses are slightly higher, we can still obtain the required beam inside the cyclotron, albeit requiring about 20% higher beam from the source.
For the IBSimu simulations a 12 mA beam composed of 80% H+${}_{2}^{+}$, 10% H+, and 10% H+${}_{3}^{+}$ was used. This provides ample current for IsoDAR assuming 50% global transmission efficiency while using current from the ion source that has already been empirically demonstrated. As illustrated in Fig. 2.9 and in Table 2.2 the LEBT can match the required input parameters for the RFQ. Moreover, as illustrated in Fig. 2.10 this match is achieved with emittance that surpasses that used for baseline RFQ simulations.
Figure 2.9: A temperature plot showing the Twiss parameters required for optimal transmission through the RFQ. The indicated point shows the parameters at the end of the LEBT. [waites_high_2022]
Figure 2.10: Phase space from the LEBT simulation (left) vs phase space for ideal beam input to the RFQ (right). From Ref. [waites_high_2022].
@@
2.2.4 Risks and Mitigations
Risk: Unexpected Power Loss. In the case of an unexpected loss of power to one of the power supplies, the beam dynamics could be very different from nominal and lead to damage downstream (e.g., the RFQ).
Mitigation: Careful simulation of various failure modes showed that the LEBT is robust under such changes. The time constant involved with heating up of materials are such that the beam can be switched off before damage occurs. However, for the MIST-2 design, we are envisioning the addition of a Faraday cup directly downstream of the extraction system. Operating spring-loaded, it can be inserted quickly to intercept the beam early on.
Risk: Lower than anticipated space charge compensation. Optimal beam transport was calculated for a high compensation factor. If this is not the case, transport efficiency and beam dynamics may deteriorate.
Mitigation: The LEBT design was double-checked with no space charge compensation and care was taken that injection into the RFQ is nevertheless possible. Typically, higher losses, which are still manageable, are seen. The ion source must provide about 20% more beam current in this case.
@@ 2.3 The Radiofrequency Quadrupole
Figure 2.11: Cut view of the IsoDAR RFQ. In the split-coaxial design, vanes are attached to the front and end plates rather than radially to the mantle. From Ref. [waites_high_2022].
@@ 2.3.1 Technical Design
The technical design was presented in Refs. [ratzingerTechnicalDesignReport2020, winklehner_high-current_2021, holtermann_technical_2021] and the optimization of the RF loop coupler in Ref. [sangroula_design_2021]. The thermal properties and cooling considerations were presented in Ref. [koser_thermal_2021]. An optimization scheme using machine learning was presented in Ref. [koser_input_2022].
Figure 2.12: Modes of the mechanical deformation of the RFQ vanes. The frequencies are 18 Hz and 25 Hz. From Ref. [holtermann_technical_2021].
@@ 2.3.2 Beam Dynamics
Beam dynamics studies were performed of the full 3D particle distribution resulting from the LEBT simulations going through the RFQ, including the effects of the entrance gap and asymmetry (due to the split coaxial design) and the exit gap and asymmetry. This was performed by the subcontracted company Bevatech GmbH in Germany, who is currently also manufacturing the RFQ. The well-established tracking code TraceWin was used [tracewin]. Fine tuning of the phase of the exit gap (i.e. tuning the length of a drift preceding the exit gap), leads to final output beam parameters listed in Tab. 2.4.
@@
Phase width (1-RMS) 40
2.3.3 Risks and Mitigation
Risk: New RFQ Technology. Direct axial injection into a compact cyclotron has never been demonstrated. It is possible that there are limitations that have not been thought of, that will prove to be show-stoppers.
Mitigation: A prototype machine is currently being built and will be thoroughly tested in the coming year. As a fallback, conventional LEBTs using a multi-harmonic buncher are well understood and reliable. However, they are much less efficient than RFQs. The focus in this case will shift to increasing ion source performance further to allow for higher beam losses during injection.
Risk: RFQ Power Estimate The RFQ costs depend strongly on the RF power necessary to drive the rods/vanes, the length of the RFQ, the engineering complexity involved in machining the vanes/rods and designing the vacuum vessel.
Mitigation: We have worked with RF engineers with intimate knowledge of RFQ design. The upcoming prototype will demonstrate the power requirements.
Risk: Beam spread between RFQ exit and spiral inflector. The simulations showed that the beam spreads transversally and longitudinally after exiting the RFQ. Reality might prove more unforgiving than computer simulations and we may incur higher beam losses than anticipated.
Mitigation: Our high-fidelity simulations suggest that the beam spread is manageable. The prototype will allow measuring this precisely. Additional focusing elements can be inserted between RFQ and spiral inflector if necessary. More aggressive collimation is possible as well, at the expense of requiring higher current from the source.
@@ 2.3.3 Risks and Mitigation
Figure 3.1: Isometric view of the cyclotron.
In this section, we describe the details of the cyclotron at the level of a preliminary design report. This section includes details of the design of the cyclotron, the spiral inflector that transmits ions from the RFQ into the cyclotron, and the extraction system.
This section also characterizes the latest in simulo performance of these systems.
@@ 3.1 Magnet Yoke
The magnet design presented here is physically realistic, but is necessarily preliminary. We do not anticipate the broad design of the magnet to change in future. As we finalize the details of beam extraction small modifications to the magnets, especially at the fringes, may accordingly be modified.
The engineering design of the IsoDAR magnet benefits from the experience gained by IBA in the fabrication of its C400 cyclotron (seen in Fig 3.2), which is slightly larger, at 7 meters outer diameter.
Figure 3.2: C400 on the milling machine, C230 in front of it. Courtesy of IBA.
@@ 3.1.1 Design and Fabrication of the Cyclotron Magnet
As with the C400, the IsoDAR cyclotron is divided into smaller parts to facilitate transport to the site. The magnet is split radially and will be assembled on site. The vacuum chamber is split into 3 self-centering parts along its height. The vacuum tightness is ensured by 2 axial O-rings at each interface except for the interface between the upper and lower chambers where 2 radial O-rings are used. An exploded visualization of the upper and lower parts is shown in Figs. 3.4 and 3.4. The material used for the magnet parts is cyclotron steel (the exact composition shall be determined after a detailed physics study), the vacuum chambers will be made of stainless steel.
@@
An cut-section view of the cyclotron is shown in Fig. 3.5. The external and vacuum disks are connected to each other as well as to the flux return using specific bolts (such as the Superbolt from Nord-Lock). The poles are split into several parts and an easily removable part should be added in the next iteration to ease the mapping and possibly improve the magnetic performance. A detailed view of the cyclotron components is shown in Fig. 3.6 and additional details about the yoke parts can be found in Table 3.1
@@
3.2 Cyclotron magnetic performance
In this section we describe the latest modifications applied to the cyclotron regarding its magnetic properties. We will compare the results with the model from the Conceptual Design Report (CDR) from [abs_isodarkamland_2015] to ensure its magnetic performance remains within specifications.
With the latest design of the vacuum chamber, see section 3.5, the pole radius was slightly reduced. To keep the magnetic field isochronous up to extraction at 120 MeV, the outer radius required small modifications, see left image in Fig. 3.7. First, the curvature was modified to follow the curvature of the last orbit. This enhances the total flux in the region where needed. Second, the top layer of the pole at its outer radius will be made of permendur. This alloy has better magnetic properties at high field strengths. Using permendur will enhance the magnetic flux. Finally, the pole height is increased by a slanted ridge at its extremity. This reduction of the pole gap should have no effect on the transmission and can be tuned to improve the dB/dr for the last orbits.
The steel used for the cyclotron yoke and poles is the SNVA6342, also used, for example, in the C400 cyclotron (see Ref. [bib:c400_iron]). Its BH-curve is shown in Fig.3.7. Here, we also see the better performance of the permendur alloy.
3.2.1 Isochronization
3.2.2 Risks and Mitigation
Risk: Manufacturing Imperfections. Because no manufacturing process is perfect, we must measure the local magnetic field developed by the accelerator magnet in an iterative process. During the mapping iterations, physicists compare measurements results with expected calculated (theoretical) magnetic field map values and propose local magnet iron shape adjustments, typically on pole tips. Difficulty of the process is a risk.
Mitigation: For that reason, pole tips, or at least the edges of the pole tips (boundary valley/hill) must be dismountable to allow re-machining and multiple tuning iterations. This mapping process is repeated until the final expected map is obtained, allowing, theoretically, the appropriate beam acceleration and extraction. In the case of IsoDAR, we can either perform the entire process underground, or the machine must be dismounted after surface tests and reinstalled underground. This will require re-measuring the magnetic field after assembly on site, during a second (shorter) mapping campaign.
Risk: Mapping Tool. Developing a mapping tool requires care as this tool accuracy will define performance accuracy of the measured magnet itself.
Mitigation: Many precautions are mandatory concerning calibrations, mechanical stiffness, thermal stability and tracking as well as the use of perfect ”non-magnetic” materials (to avoid measures perturbations and unexpected distortions).
@@ 3.3 Interface with the RFQ
Figure 3.10: RFQ overview.
The RFQ will sit on top of the cyclotron along its central axis, as shown in Fig 3.10. This places the RFQ outside of the cyclotron’s vacuum. Therefore, a vacuum sleeve will be used to ensure the integrity of the cyclotron’s vacuum. This sleeve will be fixed on top of the upper external disk and the vacuum sealing will be done radially on the upper internal disk (as seen in figure number 3.12 ). The O-rings willhave a cross sectional diameter of 8mm and an inner diameter of 305mm.
In order to allow for adjustment, the RFQ sleeve allows a rotation of 5mrad around the lower O-ring contact patch (discussed with the manufacturer of the RFQ). The RFQ sleeve is as thin as possible to avoid any loss in magnetic field in the central region (hence the O-rings in the Yoke parts).
Figure 3.11: RFQ zoom.
Figure 3.12: RFQ cut section 1.
As shown in Fig. 3.13, vacuum sealing will occur at the base of the RFQ using a thick O-ring (diameter 10 mm) which will self-seal under the RFQ’s weight ($\approx 3000$ N). The dimensions of the vacuum sleeve are listed in Table 3.2.
In order to allow fine-tuning the cyclotron alignment, a set of adjustment feet, shown in Fig. 3.11, are fixed to the cyclotron yoke. By using four groups of set screws and one vertical adjustment nut, the following alignment modifications can be made.
  • Rotation: ±$\pm 5$
@@ 3.4 Central Region
Figure 3.14: CAD rendering of the spiral inflector and particle trajectories as implemented previously for early tests at Best Cyclotron Systems, Inc. [alonso_isodar_2015]. The positive and negative electrostatic deflectors are shown in blue and green respectively. The cyclotron magnetic field in this image is directed vertically upwards. Particles enter the spiral inflector via the rectangular grounded collimator (solid gray) and are guided into the cyclotron mid-plane by means of the cyclotron magnetic field and the electrostatic potential between the electrodes. The copper housing (transparent gray) isolates the spiral inflector from the RF fields driving the cyclotron. From Ref. [winklehner_realistic_2017].
@@
The injection system design is a crucial component that allows us to achieve a high beam current. Our design balances injecting the $\mathrm{H}_{2}^{+}$ molecule at high energy against avoiding the use of complex high-voltage platforms. Higher energy injection mitigates space charge and reduces geometrical (non-normalized) emittance. An energy of 35 keV/amu was selected. This in turn determines the target output energy of the RFQ. As in commercial cyclotrons, an axial injection system based on a Spiral Inflector (SI) will be used to bend the beam from the axial direction to the median plane. We have investigated several refinements to a traditional spiral inflector which will be discussed subsequently, but the concept, as implemented successfully in a previous iteration of this work and tested at Best Cyclotron Systems, Inc., is visualized in Figure 3.14 using trajectories calculated with OPAL [adelmann_opal_2019].
@@
The exit point of the spiral inflector is also the starting point of acceleration in the Central Region (CR) of the cyclotron. First turn acceleration is then achieved by the dee tips extending into recesses in the spiral inflector housing. The shapes of the dee tips and the spiral inflector housing have been designed to guide the particles from the spiral inflector exit to the acceleration region while providing the necessary energy gain and beam focusing.
The central region and the spiral inflector must be designed carefully and together because they have strong interplay. This procedure has already undergone one past iteration in the summers of 2013 and 2014 at Best Cyclotron Systems Inc. In the following sections, we will discuss improvements to this design which will be incorporated into the preliminary design for the full IsoDAR machine.
3.4.1 Spiral Inflector Overview & Design Considerations
The axial injection of the ionized beam into a cyclotron is realized using an electrostatic device called a spiral inflector, which consists of two curved electrode deflectors (see Figure 3.14). The electrostatic potential between these electrodes is able to bend the beam 90 from the axial line to the median plane of the cyclotron. The helical trajectory of the beam is determined both by the shape of the electrodes, the electrode potentials, and by the magnetic field of the cyclotron as the beam is bent into the median plane. In order to isolate the inflector region from external fields as much as possible, grounded entrance and exit apertures and a grounded copper housing are also included.
An ideal spiral inflector would do nothing more than rotate the beam by 90, introducing no additional beam divergence or losses and seamlessly merging with initial accelerating in the central region. Further, in order to achieve this degree of bending in the small space of the central region, $\mathcal{O}(kV)$ potentials are required, so it is important to minimize or even eliminate any $\mathrm{H}_{2}^{+}$ impact on the electrodes to prevent sparking.
It is also important to consider the impact of a realistic magnetic field when designing our spiral inflector. This is particularly important for the IsoDAR inflector because of its relatively large size. Unlike other spiral inflectors that have been constructed, we will require a larger gap and plate size due to the non-negligible space charge and higher rigidity of $\mathrm{H}_{2}^{+}$. This larger volume means that variations in the magnetic field along the trajectory are non-negligible. Accordingly, all following work uses a full 3D magnetic field map generated using OPERA.
3.4.2 Designing the Inflector
Ultimately, an additional iteration of production and empirical testing of the final IsoDAR inflector will be performed. To optimize the inflector beforehand, we use a numerical optimization procedure to inform the best theoretical parameters.
For an inflector in a uniform magnetic field, analytic solutions to particle tracking exist for the central trajectory. However when 3D magnetic variations and/or off center trajectories are considered a numerical approach is required.
An initial estimate based on an analytical solution for a central track serves as a numerical starting point, after which the mesh generating software gmsh [GMSHarticle] is used to produce surface representations which inform subsequent electrostatic analysis. Python’s BEMPP API [bempp] is then used to perform boundary analysis.
In short, this procedure enables a geometric model of our inflector to yield a full 3D electrostatic. Particles can then be tracked with high precision using the Boris method [boris].
A major strength of this procedure and choice of software is speed, with a full simulation iteration taking only $\mathcal{O}$(hours). Prior work [weigelspiral] comparing this process to commercial software such as COMSOL demonstrates that the differences in field calculations are negligible.
3.4.3 Geometric Refinements
In most spiral inflectors, the two electrodes would be rectangular prisms if flattened out. They remain parallel to each other with locally flat surfaces while spiraling. Hereafter we will call this a “Traditional” spiral inflector. A traditional inflector can effectively transfer particles, but for off center paths these inflectors can introduce beam divergences in energy and position which were not present in the initial beam.
We address this by adding several refinements to prior inflector iterations.
  • Beam focusing between the RFQ and spiral inflector
  • Introducing V-shape
  • Introducing plate angling
  • Modifying the terminal end to a wedge
The addition of intermittant beam focusing ensures that the $\mathrm{H}_{2}^{+}$ entering the spiral inflector are as near to the central trajectory as possible. This is accomplished by the addition of electrostatic quadrupoles in the gap between the RFQ and inflector entrance (see also Fig. 1.7). Using the output of the RFQ simulation [rfqsim] as input to the quadrupoles, geometry and potentials were optimized such that the tightest possible input beam is produced.
The remaining three modifications: V-shape, plate angling, and the terminal wedge were motivated by work done by Barnard et al. [barnard_inflector]. These all have the effect of introducing quadrupole moments to the field of the inflector itself, which can provide mid-track focusing to minimize beam dispersion in flight.
The V-shape introduces a concave v-shape to the anode and a convex v-shape to the cathode when viewed along the long axis of the electrode (Fig.3.15(a)). Plate angling reorients the plate angle relativce to each other along the course of the inflector. In other words the plates are parallel only in the center with a positive angle between the plates initially linearly transitioning to an equal and opposite negative angle by the end (Fig.3.15(b)). The terminal gamma wedge cut modifies the end of the anode into a triangular rather than flat profile at the end and shortens the cathode. This assists with the fringe fields at the end which can produce vertical spreading. (Fig.3.15(c)).
(a) V-Shape Illustration
(b) Angling Illustration
(c) Gamma Wedge Illustration
Figure 3.15: Illustrations of different focusing concepts. From Ref. [barnard_inflector].
These modifications and their associated focusing fields interplay, but can be independently varied alongside other parameters like electrode potential, electrode separation, etc. We are actively exploring different techniques to handle this high-dimensional optimization problem. Monte Carlo sampling is typically well suited to this sort of optimization, but studies in which only one parameter is varied at a time have illustrated that optimization extrema can be relatively narrow, necessitating a significant, if not intractable, amount of computation time.
As we are at the edge of practical optimization using Monte Carlo, this leads us to consider other potential alternatives. Surrogate models using a fully connected neural network capable of learning and interpolating underlying features have been previously explored[EdelenMachingLearning] [KoserMachineLearning] [villarrealMachineLearning]. Initial iterations have proved promising and have informed the parameters for our current best inflector performance. As this work matures, it will be detailed in upcoming publications.
3.4.4 Characterization of Spiral Inflector
The result of adding optimized focusing quadrupoles on the beam distribution entering the spiral inflector is illustrated in Fig. 3.16.
Figure 3.16: Illustration of beam tightness at the entrance of the spiral inflector before and after the addition of focusing quadrupoles between the RFQ and inflector
The optimized spiral inflector, with $\mathrm{H}_{2}^{+}$ tagged through, is illustrated in Fig.3.17.
Figure 3.17: Illustrations of the optimized spiral inflector. Note the v-shape to the electrodes as well as the non-parallel angle seen in the bottom view, and the angled wedge visible in the top view.
The efficiency of this modified inflector is significantly improved. Efficiency must be thought of from two fronts. On the one hand we wish to transmit as much current from the RFQ into the cyclotron as possible.Particles which are not successfully transmitted must terminate somewhere. Termination on grounded apertures is acceptable while termination on the HV electrodes will result in sparking and can have significantly impact on running. It is therefore also vital to specifically consider the fraction of ions that terminate on an electrode.
We thus consider two similar cases. Both have identical optimized spiral inflectors. The difference is whether or not a grounded collimator is placed at the entrance to further narrow the entering beam beyond what the quadrupoles have provided. This collimator is a circular aperture which, for this study, was chosen to be as large as possible while guaranteeing that less than one percent of $\mathrm{H}_{2}^{+}$ impact the electrodes. The resulting efficiencies and $\mathrm{H}_{2}^{+}$-electrode collisions are given in Table 3.3.
Table 3.3: Comparison of efficiency and ion-electrode impact for the optimized vs traditional spiral inflector.
Efficiency Electrode Impact %
Traditional Inflector 74% 17%
Optimized Inflector 97% 1%
Traditional Inflector (Coll.) 39% $<1$%
Optimized Inflector (Coll.) 93% $<1$%
We also need to consider the impact of the new inflector on the quality of the beam which is successfully transmitted. These changes are characterized in Table 3.4. Detailed histograms showing the kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ at the entrance and exit of the inflector are shown in Fig. 3.18.
Table 3.4: Comparison of transmitted beam parameters at the inflector exit.
$\sigma_{x}$ [mm] $\sigma_{y}$ [mm] $\sigma_{z}$ [mm] $\sigma_{r}$ [mm] $\sigma_{KE}$ [keV]
Traditional Inflector 3.0 3.9 10.1 2.3 4.0
Optimized Inflector 4.9 2.7 6.4 2.4 3.1
Figure 3.18: Illustration of the kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ prior to the inflector entrance vs at the inflector exit.
3.4.5 Orbit Matching and Beam Centering
Figure 3.19: Beam tracking result from the AIMA central region study. Left beam centering and acceleration on top of a view of the electrodes. Colors correspond to prticle phase. Right: 1 cm turn separation in the fourth turn with a single collimator. From [aima:central].
Once the beam emerges from the spiral inflector, it must be properly matched in terms of orbit radius, radial momentum, and phase. A specially designed shape of the inner ends of the RF cavities performs this duty. The angles and widths of the accelerating gaps of each cavity are modified in an optimization procedure during the first two turns to place the beam onto a static equilibrium orbit appropriate for the respective energy. The center of a full integrated orbit after these two turns should then align well with the cyclotron center.
The spiral inflector end can be tilted (often this parameter is called $k^{\prime}$) to aid in the matching and thus becomes part of the optimization procedure.
A preliminary design study of the IsoDAR central region was performed by the company AIMA in France [aima:central]. The AIMA central region model is shown in Fig. 3.19. This central region was designed with a slightly different magnetic pole shape and with its own matched spiral inflector, using the ouput beam from an early RFQ beam tracking and single electrostatic quadrupole transverse refocusing element. It demonstrated that the RFQ-generated beam could be injected and vertically focused. By using a single collimator (removing $\approx 42$% of the beam), an edge-to-edge turn separation of 1 cm was achieved in the fourth turn (approximately 1 MeV/amu).
The RMS beam size and energy spread was significantly worse than what we see in our optimized inflector. Furthermore, the AIMA study did not consider space charge, which is fundamental to establishing vortex motion (see below). Somewhat counterintuitively, including space-charge will improve the results further; as will our new optimized spiral inflector.
In the next step, we will modify the central region and optimized spiral inflector to match well and track particles through the finalized geometry with OPAL using the PIC method to demonstrate a stable vortex.
For this preliminary design, we use a Gaussian beam with size and emittance informed by the AIMA study for further simulations of acceleration to 60 MeV/amu.
Figure 3.20: Reduction of beam halo by collimators
3.4.6 Low Energy Collimation
Following injection, there is significant beam spread in the first few turns of the cyclotron that we need to manage. We do this by placing a series of collimators around the beam core in the early turns. We must do all of our collimation in the first 7 turns of the primary acceleration, corresponding approximately to a 2 MeV/amu limit, to avoid activating our collimators. The placement of the collimators is shown in Figure  3.21 The reduction in RMS spread is shown in Fig  3.41. In total, we have 15 collimators that remove 36.1% of the beam. The improvement in beam shape is shown in  3.20. With time, the halo slowly redevelops but to a lesser extent than it would without collimation.
Figure 3.21: Placement of 15 collimators in first 7 turns
It is necessary to make the beam as circular as possible following the low energy collimators so that the beam is well-matched to “vortex motion.” Vortex motion stabilizes the beam and significantly reduces beam growth over time. It arises from the interaction between the self-field of the beam and the external magnetic field of the cyclotron. Vortex motion is discussed in depth in Ref [baumgarten:vortex2].
3.4.7 Risks and Mitigation
Risk: High voltage on the spiral inflector electrodes. The high voltage on the spiral inflector electrodes can lead to arcing from one electrode to the other or to ground, creating unstable injection conditions.
Mitigation: During the BCS tests we were able to run the spiral inflector electrodes as high as ±13.5 kV without beam using conventional polishing and cleaning methods for the electrodes. Electro-polishing and conditioning in an inert gas environment will improve this even more. Beam striking the electrodes could potentially induce arcs through electron emission. This can be avoided by presenting a well-defined beam to the spiral inflector to assure close to 100% transmission. We have and will continue to emphasize minimizing or completely eliminating ion impacts in simulo with these tools. Finally, we will set up machine protection algorithms that can break a spark and restart acceleration in a controlled manner to minimize downtimes due to sparking.